OpenStudy (anonymous):

jacks father bough a used boat he paided $6,937.89 for the boat and $1,843.56 for extras what was the total cost of the boat

OpenStudy (anonymous):

6937.89 + 1843.56 = 8781.45

OpenStudy (anonymous):

(6937.89+1843.56)= 175629/20 or 8781.45